# The Sample-Communication Complexity Trade-off in Federated Q-Learning

### v1

@@ 2.3 Intermittent-communication algorithm protocols
We consider a popular class of federated learning algorithms with intermittent communication. The intermittent communication setting provides a natural framework to extend single-agent Q-learning algorithms to the distributed setting. As the name suggests, under this setting, the agents intermittently communicate with each other or a central server, sharing their updated beliefs about $Q^{\star}$. Between two communication rounds, each agent updates their belief about $Q^{\star}$ using stochastic approximation iterations based on the locally available data, similar to a single-agent setup. Such intermittent communication algorithms have been extensively studied and used to establish lower bounds on communication complexity of distributed stochastic convex optimization (Woodworth et al., 2018, 2021).
@@ 3 Communication Complexity Lower Bound
The above lower bound on the communication complexity of federated Q-learning is a consequence of the bias-variance trade-off that governs the convergence of the Q-learning algorithm. While a careful choice of step sizes alone is sufficient to balance this trade-off in the centralized setting, the choice of communication schedule also plays an important role in balancing this trade-off in the federated setting. The local steps between two communication rounds induce a positive estimation bias that depends on the standard deviation of the iterates and is a well-documented issue of “over-estimation” in Q-learning (Hasselt, 2010). Since such a bias is driven by local updates, it does not reflect any benefit of collaboration. During a communication round, the averaging of iterates across agents allows the algorithm an opportunity to counter this bias by reducing the effective variance of the updates through averaging. In our analysis, we show that if the communication is infrequent, the local bias becomes the dominant term and averaging of iterates is insufficient to counter the impact of the positive bias induced by the local steps. As a result, we do not observe any statistical gains when the communication is infrequent. Our argument is inspired by the analysis of Q-learning in Li et al. (2024) and is based on analyzing the convergence of an intermittent communication algorithm on a specifically chosen “hard” MDP instance. The detailed proof is deferred to Appendix A.
@@
Several recent studies (Liu and Olshevsky, 2023; Tian et al., 2024) established that a single round of communication is sufficient to achieve linear speedup of TD learning for policy evaluation, which do not contradict with our results focusing on Q-learning for policy learning. The latter is more involved due to the nonlinearity of the Bellman optimality operator.
@@
We would like to point out that our lower bound extends to the asynchronous setting (Li et al., 2024) as the assumption of i.i.d. noise corresponding to a generative model is a special case of Markovian noise observed in the asynchronous setting.
@@ 4 The Fed-DVR-Q Algorithm
We propose a new federated Q-learning algorithm, Fed-DVR-Q, that achieves not only a communication complexity of $\textsf{CC}_{\textsf{round}}=\widetilde{\mathcal{O}}\left(\frac{1}{1-\gamma}\right)$ and $\textsf{CC}_{\textsf{bit}}=\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\right)$ but also order-optimal sample complexity (up to logarithmic factors), thereby providing a tight characterization of the achievability frontier that matches with the converse result derived in the previous section.
@@ 4.1 Algorithm description
1:  Input : Error bound $\varepsilon>0$, failure probability $\delta>0$
2:  $k\leftarrow 1,Q^{(0)}\leftarrow\mathbf{0}$
3:  // Set parameters as described in Section 4.1.3
4:  for $k=1,2,\dots,K$ do
5:     $Q^{(k)}\leftarrow\textsc{RefineEstimate}(Q^{(k-1)},B,I,{L}_{k},D_{k},J)$
6:     $k\leftarrow k+1$
7:  end for
8:  return $Q^{(K)}$
Algorithm 2 Fed-DVR-Q
@@ 4.1.1 The RefineEstimate sub-routine
RefineEstimate, starting from $\overline{Q}$, an initial estimate of $Q^{\star}$, uses variance-reduced Q-learning updates (Wainwright, 2019b; Sidford et al., 2018) to obtain an improved estimate of $Q^{\star}$. It is characterized by four parameters — the initial estimate $\overline{Q}$, the number of local iterations $I$, the re-centering sample size $L$ and the batch size $B$, which can be appropriately tuned to control the quality of the returned estimate. Additionally, it also takes input two parameters $D$ and $J$ required by the compressor $\mathscr{C}(\cdot;D,J)$, whose description is deferred to Section 4.1.2.
@@ 4.1.3 Setting the parameters
where $K_{0}=\lceil\frac{1}{2}\log_{2}(\frac{1}{1-\gamma})\rceil$. The piecewise definition of $L_{k}$ is crucial to obtain the optimal dependence with respect to $\frac{1}{1-\gamma}$, similar to the two-step procedure outlined in Wainwright (2019b).
@@ 4.2 Performance guarantees
A proof of Theorem 2 can be found in Appendix B. A few implications of the theorem are in order.
@@
The commonly adopted approach in intermittent communication algorithm is to use a local update scheme that takes multiple small (i.e., $B=\mathcal{O}(1)$), noisy updates between communication rounds, as evident from the algorithm design in Khodadadian et al. (2022); Woo et al. (2023) and even numerous FL algorithms for stochastic optimization (McMahan et al., 2017; Haddadpour et al., 2019; Khaled et al., 2020). In Fed-DVR-Q, we replace the local update scheme of taking multiple small, noisy updates by a single, large update with smaller variance, obtained by averaging the noisy updates over a minibatch of samples. The use of updates with smaller variance in variance-reduced Q-learning yields the algorithm its name. While both the approaches result in similar sample complexity guarantees, the local update scheme requires more frequent averaging across clients to ensure that the bias of the estimate, also commonly referred to as “client drift”, is not too large. On the other hand, the minibatching approach does not encounter the problem of bias accumulation from local updates and hence can afford more infrequent averaging, allowing Fed-DVR-Q to achieve optimal order of communication complexity.
@@
Fed-DVR-Q is the first federated Q-learning algorithm to analyze and establish communication complexity at the bit level. While all existing studies on federated Q-learning focus only on the frequency of communication and assume transmission of real numbers with infinite bit precision, our analysis provides a more holistic view point of communication complexity at the bit level, which is of great practical significance. While some recent other studies (Wang et al., 2023) also considered quantization in federated RL, their objective is to understand the impact of message size on the convergence with no constraint on the frequency of communication, unlike the holistic viewpoint adopted in this work.
@@ 5 Conclusion
We presented a complete characterization of the sample-communication trade-off for federated Q-learning algorithms with intermittent communication. We showed that no federated Q-learning algorithm with intermittent communication can achieve any speedup with respect to the number of agents if its number of communication rounds are sublinear in $\frac{1}{1-\gamma}$. We also proposed a new federated Q-learning algorithm called Fed-DVR-Q that uses variance reduction along with minibatching to achieve optimal-order sample and communication complexities. In particular, we showed that Fed-DVR-Q has a per-agent sample complexity of $\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{M(1-\gamma)^{3}\varepsilon^{2}}\right)$, which is order-optimal in all salient problem parameters, and a communication complexity of $\widetilde{\mathcal{O}}\left(\frac{1}{1-\gamma}\right)$ rounds and $\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\right)$ bits.
@@
Acknowledgement
This work is supported in part by the grants NSF CCF-2007911, CCF-2106778, CNS-2148212, ECCS-2318441, ONR N00014-19-1-2404 and AFRL FA8750-20-2-0504, and in part by funds from federal agency and industry partners as specified in the Resilient & Intelligent NextG Systems (RINGS) program.
@@ Appendix A Proof of Theorem 1
The proof builds upon analyzing the behavior of a generic algorithm $\mathscr{A}$ outlined in Algorithm 1 over a particular instance of MDP. The particular choice of MDP is inspired by, and borrowed from, other lower bound proofs in the single-agent setting [Li et al., 2024] and helps highlight core issues that lie at the heart of the sample-communication complexity trade-off. Following Li et al. [2024], the construction is first over a small state-action space that allows us to focus on a simpler problem before generalizing it to larger state-action spaces.
@@
In the second step, we analyze the error of the iterates generated by an intermittent communication algorithm $\mathscr{A}$. The analysis is inspired by the single-agent analysis in Li et al. [2024], which highlights the underlying bias-variance trade-off. Through careful analysis of the algorithm dynamics in the federated setting, we uncover the impact of communication on the bias-variance trade-off and the resulting error of the iterates to obtain the lower bound on the communication complexity.
@@ A.1 Introducing the “hard” instance
We first introduce an MDP instance $\mathcal{M}_{h}$ that we will use throughout the proof to establish the result. Note that this MDP is identical to the one considered in Li et al. [2024] to establish the lower bounds on the performance of single-agent Q-learning algorithm. It consists of four states $\mathcal{S}=\{0,1,2,3\}$. Let $\mathcal{A}_{s}$ denote the action set associated with the state $s$. The probability transition kernel and the reward function of $\mathcal{M}_{h}$ is given as follows:
@@ A.3.1 Small learning rates
In this subsection, we prove the lower bound for small learning rates, which follow from similar arguments in Li et al. [2024].
For this case, we focus on the dynamics of state $2$. We claim that the same relation established in Li et al. [2024] continues to hold, which will be established momentarily:
@@
To obtain lower bound on $V^{\star}(2)-\mathbb{E}[V_{T}^{m}(2)]$, we need to obtain a lower bound on $\eta_{0}^{(T)}$, which from [Li et al., 2024, Eqn. (120)] we have
@@ A.3.3 Large learning rates with large $\frac{\eta_{T}}{BM}$
In order to bound the error in this scenario, note that $\frac{\eta_{T}}{BM}$ controls the variance of the stochastic updates in the fixed point iteration. Thus, when $\frac{\eta_{T}}{BM}$ is large, the variance of the iterates is large, resulting in a large error. To demonstrate this effect, we focus on the dynamics of state $2$. This part of the proof is similar to the large learning rate case of Li et al. [2024]. For all $t\in[T]$, define:
@@
    The analysis for this case is identical to that considered in Li et al. [2024]. We explicitly write the steps for completeness. Specifically,
@@
    Using [Li et al., 2024, Eqn.(134)], for any $t^{\prime}<t$ and all agents $m$, we have the relation
    $V_{t}^{m}(2)=\frac{1}{1-\gamma p}-\prod_{k=t^{\prime}+1}^{t}(1-\eta_{k}(1-\gamma p))\left(\frac{1}{1-\gamma p}-V_{t^{\prime}}^{m}(2)\right)+\sum_{k=t^{\prime}+1}\eta_{k}^{(t)}\gamma(\widehat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2).$
@@
    $\overline{V}_{T}(2)=\frac{1}{1-\gamma p}-\prod_{k=t^{\prime}+1}^{T}(1-\eta_{k}(1-\gamma p))\left(\frac{1}{1-\gamma p}-\overline{V}_{t^{\prime}}(2)\right)+\frac{1}{M}\sum_{m=1}^{M}\sum_{k=t^{\prime}+1}^{T}\eta_{k}^{(T)}\gamma(\widehat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2).$ (83)
    Let $\{\mathscr{F}_{t}\}_{t=0}^{T}$ be a filtration such that $\mathscr{F}_{t}$ is the $\sigma$-algebra corresponding to $\{\{\widehat{P}^{m}_{s}(2|2)\}_{m=1}^{M}\}_{s=1}^{t}$. It is straightforward to note that $\left\{\frac{1}{M}\sum_{m=1}^{M}\eta_{k}^{(T)}\gamma(\widehat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\right\}_{k}$ is a martingale sequence adapted to the filtration $\mathscr{F}_{k}$. Thus, using the result from [Li et al., 2024, Eqn.(139)], we can conclude that
    $\text{Var}(\overline{V}_{T}(2))\geq\mathbb{E}\left[\sum_{k=\tau^{\prime}+2}^{T}\text{Var}\left(\frac{1}{M}\sum_{m=1}^{M}\eta_{k}^{(T)}\gamma(\widehat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\ \bigg{|}\mathscr{F}_{k-1}\right)\right].$ (84)
@@
    $\text{Var}\left(\frac{1}{M}\sum_{m=1}^{M}\eta_{k}^{(T)}\gamma(\widehat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\ \bigg{|}\mathscr{F}_{k-1}\right)$ $=\frac{1}{M^{2}}\sum_{m=1}^{M}\text{Var}\left(\eta_{k}^{(T)}\gamma(\widehat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\ \bigg{|}\mathscr{F}_{k-1}\right)$

### v2

@@ 2.3 Intermittent-communication algorithm protocols
We consider a popular class of federated learning algorithms with intermittent communication. The intermittent communication setting provides a natural framework to extend single-agent Q-learning algorithms to the distributed setting. As the name suggests, under this setting, the agents intermittently communicate with each other or a central server, sharing their updated beliefs about $Q^{\star}$. Between two communication rounds, each agent updates their belief about $Q^{\star}$ using stochastic approximation iterations based on the locally available data, similar to a single-agent setup. Such intermittent communication algorithms have been extensively studied and used to establish lower bounds on communication complexity of distributed stochastic convex optimization (Woodworth2018GraphOracle; Woodworth2021IntermittentComm).
@@ 3 Communication Complexity Lower Bound
The above lower bound on the communication complexity of federated Q-learning is a consequence of the bias-variance trade-off that governs the convergence of the Q-learning algorithm. While a careful choice of step sizes alone is sufficient to balance this trade-off in the centralized setting, the choice of communication schedule also plays an important role in balancing this trade-off in the federated setting. The local steps between two communication rounds induce a positive estimation bias that depends on the standard deviation of the iterates and is a well-documented issue of “over-estimation” in Q-learning (Hasselt2010DoubleQL). Since such a bias is driven by local updates, it does not reflect any benefit of collaboration. During a communication round, the averaging of iterates across agents allows the algorithm an opportunity to counter this bias by reducing the effective variance of the updates through averaging. In our analysis, we show that if the communication is infrequent, the local bias becomes the dominant term and averaging of iterates is insufficient to counter the impact of the positive bias induced by the local steps. As a result, we do not observe any statistical gains when the communication is infrequent. Our argument is inspired by the analysis of Q-learning in Li2023QLMinimax and is based on analyzing the convergence of an intermittent communication algorithm on a specifically chosen “hard” MDP instance. The detailed proof is deferred to Appendix A.
@@
Several recent studies (liu2023distributed; tian2024one) established that a single round of communication is sufficient to achieve linear speedup of TD learning for policy evaluation, which do not contradict with our results focusing on Q-learning for policy learning. The latter is more involved due to the nonlinearity of the Bellman optimality operator. Specifically, if the operator whose fixed point is to be found is linear in the decision variable (e.g., the value function in TD learning) then the fixed point update only induces a variance term corresponding to the noise. However, if the operator is non-linear, then in addition to the variance term, we also obtain a bias term in the fixed point update. While the variance term can be controlled with one-shot averaging, more frequent communication is necessary to ensure that the bias term is small enough.
@@
We would like to point out that our lower bound extends to the asynchronous setting (Li2023QLMinimax) as the assumption of i.i.d. noise corresponding to a generative model is a special case of Markovian noise observed in the asynchronous setting.
@@ 4 The Fed-DVR-Q  Algorithm
We propose a new federated Q-learning algorithm, Fed-DVR-Q , that achieves not only a communication complexity of $\textsf{CC}_{\textsf{round}}=\tilde{\mathcal{O}}\left(\frac{1}{1-\gamma}\right)$ and $\textsf{CC}_{\textsf{bit}}=\tilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\right)$ but also order-optimal sample complexity (up to logarithmic factors), thereby providing a tight characterization of the achievability frontier that matches with the converse result derived in the previous section.
@@ 4.1 Algorithm description
{wrapfigure}
R0.6 1:  Input : Error bound $\varepsilon>0$, failure probability $\delta>0$ 2:  $k\leftarrow 1,Q^{(0)}\leftarrow\mathbf{0}$ 3:  // Set parameters as described in Section 4.1.3 4:  for $k=1,2,\dots,K$ do 5:     $Q^{(k)}\leftarrow\textsc{RefineEstimate}(Q^{(k-1)},B,I,{L}_{k},D_{k},J)$ 6:     $k\leftarrow k+1$ 7:  end for 8:  return $Q^{(K)}$ Algorithm 2 Fed-DVR-Q
@@ 4.1.1 The RefineEstimate sub-routine
RefineEstimate, starting from $\overline{Q}$, an initial estimate of $Q^{\star}$, uses variance-reduced Q-learning updates (Wainwright2019VarianceReduced; sidford2018near) to obtain an improved estimate of $Q^{\star}$. It is characterized by four parameters — the initial estimate $\overline{Q}$, the number of local iterations $I$, the re-centering sample size $L$ and the batch size $B$, which can be appropriately tuned to control the quality of the returned estimate. Additionally, it also takes input two parameters $D$ and $J$ required by the compressor $\mathscr{C}(\cdot;D,J)$, whose description is deferred to Section 4.1.2.
@@ 4.1.3 Setting the parameters
where $K_{0}=\lceil\frac{1}{2}\log_{2}(\frac{1}{1-\gamma})\rceil$. The piecewise definition of $L_{k}$ is crucial to obtain the optimal dependence with respect to $\frac{1}{1-\gamma}$, similar to the two-step procedure outlined in Wainwright2019VarianceReduced.
@@ 4.2 Performance guarantees
A proof of Theorem 2 can be found in Appendix LABEL:appendix:dvr_analysis. A few implications of the theorem are in order.
@@
The commonly adopted approach in intermittent communication algorithm is to use a local update scheme that takes multiple small (i.e., $B=\mathcal{O}(1)$), noisy updates between communication rounds, as evident from the algorithm design in Khodadadian2022FederatedQL; Woo2023FedSynQ and even numerous FL algorithms for stochastic optimization (Mcmahan2017FedAvg; Haddadpour2019LUPASGD; Khaled2020LocalSGDHeterogenous). In Fed-DVR-Q , we replace the local update scheme of taking multiple small, noisy updates by a single, large update with smaller variance, obtained by averaging the noisy updates over a minibatch of samples. The use of updates with smaller variance in variance-reduced Q-learning yields the algorithm its name. While both the approaches result in similar sample complexity guarantees, the local update scheme requires more frequent averaging across clients to ensure that the bias of the estimate, also commonly referred to as “client drift”, is not too large. On the other hand, the minibatching approach does not encounter the problem of bias accumulation from local updates and hence can afford more infrequent averaging, allowing Fed-DVR-Q  to achieve optimal order of communication complexity.
@@
Fed-DVR-Q  is the first federated Q-learning algorithm to analyze and establish communication complexity at the bit level. While all existing studies on federated Q-learning focus only on the frequency of communication and assume transmission of real numbers with infinite bit precision, our analysis provides a more holistic view point of communication complexity at the bit level, which is of great practical significance. While some recent other studies (Wang2023FederatedTDHeterogeneity) also considered quantization in federated RL, their objective is to understand the impact of message size on the convergence with no constraint on the frequency of communication, unlike the holistic viewpoint adopted in this work.
@@ 5 Conclusion
We presented a complete characterization of the sample-communication trade-off for federated Q-learning algorithms with intermittent communication. We showed that no federated Q-learning algorithm with intermittent communication can achieve any speedup with respect to the number of agents if its number of communication rounds are sublinear in $\frac{1}{1-\gamma}$. We also proposed a new federated Q-learning algorithm called Fed-DVR-Q  that uses variance reduction along with minibatching to achieve optimal-order sample and communication complexities. In particular, we showed that Fed-DVR-Q  has a per-agent sample complexity of $\tilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{M(1-\gamma)^{3}\varepsilon^{2}}\right)$, which is order-optimal in all salient problem parameters, and a communication complexity of $\tilde{\mathcal{O}}\left(\frac{1}{1-\gamma}\right)$ rounds and $\tilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\right)$ bits.
@@ Appendix A Proof of Theorem 1
The proof builds upon analyzing the behavior of a generic algorithm $\mathscr{A}$ outlined in Algorithm 1 over a particular instance of MDP. The particular choice of MDP is inspired by, and borrowed from, other lower bound proofs in the single-agent setting (Li2023QLMinimax) and helps highlight core issues that lie at the heart of the sample-communication complexity trade-off. Following Li2023QLMinimax, the construction is first over a small state-action space that allows us to focus on a simpler problem before generalizing it to larger state-action spaces.
@@
In the second step, we analyze the error of the iterates generated by an intermittent communication algorithm $\mathscr{A}$. The analysis is inspired by the single-agent analysis in Li2023QLMinimax, which highlights the underlying bias-variance trade-off. Through careful analysis of the algorithm dynamics in the federated setting, we uncover the impact of communication on the bias-variance trade-off and the resulting error of the iterates to obtain the lower bound on the communication complexity.
@@ A.1 Introducing the “hard” instance
We first introduce an MDP instance $\mathcal{M}_{h}$ that we will use throughout the proof to establish the result. Note that this MDP is identical to the one considered in Li2023QLMinimax to establish the lower bounds on the performance of single-agent Q-learning algorithm. It consists of four states $\mathcal{S}=\{0,1,2,3\}$. Let $\mathcal{A}_{s}$ denote the action set associated with the state $s$. The probability transition kernel and the reward function of $\mathcal{M}_{h}$ is given as follows:
@@ A.3.1 Small learning rates
In this subsection, we prove the lower bound for small learning rates, which follow from similar arguments in Li2023QLMinimax.
For this case, we focus on the dynamics of state $2$. We claim that the same relation established in Li2023QLMinimax continues to hold, which will be established momentarily:
@@
To obtain lower bound on $V^{\star}(2)-\mathbb{E}[V_{T}^{m}(2)]$, we need to obtain a lower bound on $\eta_{0}^{(T)}$, which from (Li2023QLMinimax, Eqn. (120)) we have
@@ A.3.3 Large learning rates with large $\frac{\eta_{T}}{BM}$
In order to bound the error in this scenario, note that $\frac{\eta_{T}}{BM}$ controls the variance of the stochastic updates in the fixed point iteration. Thus, when $\frac{\eta_{T}}{BM}$ is large, the variance of the iterates is large, resulting in a large error. To demonstrate this effect, we focus on the dynamics of state $2$. This part of the proof is similar to the large learning rate case of Li2023QLMinimax. For all $t\in[T]$, define:
@@
    The analysis for this case is identical to that considered in Li2023QLMinimax. We explicitly write the steps for completeness. Specifically,
@@
    Using (Li2023QLMinimax, Eqn.(134)), for any $t^{\prime}<t$ and all agents $m$, we have the relation
    $V_{t}^{m}(2)=\frac{1}{1-\gamma p}-\prod_{k=t^{\prime}+1}^{t}(1-\eta_{k}(1-\gamma p))\left(\frac{1}{1-\gamma p}-V_{t^{\prime}}^{m}(2)\right)+\sum_{k=t^{\prime}+1}\eta_{k}^{(t)}\gamma(\hat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2).$
@@
    $\overline{V}_{T}(2)=\frac{1}{1-\gamma p}-\prod_{k=t^{\prime}+1}^{T}(1-\eta_{k}(1-\gamma p))\left(\frac{1}{1-\gamma p}-\overline{V}_{t^{\prime}}(2)\right)+\frac{1}{M}\sum_{m=1}^{M}\sum_{k=t^{\prime}+1}^{T}\eta_{k}^{(T)}\gamma(\hat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2).$ (83)
    Let $\{\mathscr{F}_{t}\}_{t=0}^{T}$ be a filtration such that $\mathscr{F}_{t}$ is the $\sigma$-algebra corresponding to $\{\{\hat{P}^{m}_{s}(2|2)\}_{m=1}^{M}\}_{s=1}^{t}$. It is straightforward to note that $\left\{\frac{1}{M}\sum_{m=1}^{M}\eta_{k}^{(T)}\gamma(\hat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\right\}_{k}$ is a martingale sequence adapted to the filtration $\mathscr{F}_{k}$. Thus, using the result from (Li2023QLMinimax, Eqn.(139)), we can conclude that
    $\text{Var}(\overline{V}_{T}(2))\geq\mathbb{E}\left[\sum_{k=\tau^{\prime}+2}^{T}\text{Var}\left(\frac{1}{M}\sum_{m=1}^{M}\eta_{k}^{(T)}\gamma(\hat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\ \bigg{|}\mathscr{F}_{k-1}\right)\right].$ (84)
@@
    $\text{Var}\left(\frac{1}{M}\sum_{m=1}^{M}\eta_{k}^{(T)}\gamma(\hat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\ \bigg{|}\mathscr{F}_{k-1}\right)$ $=\frac{1}{M^{2}}\sum_{m=1}^{M}\text{Var}\left(\eta_{k}^{(T)}\gamma(\hat{P}_{k}^{m}(2|2)-p)V_{k-1}^{m}(2)\ \bigg{|}\mathscr{F}_{k-1}\right)$